OpenStudy (anonymous):

Jesse rents a moving van for $75 and must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount and drove the same number of miles, how far did they each travel? I got 10 miles, could anyone check that for me?

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

10 miles yeh

OpenStudy (anonymous):

I used 75+2x=80+1.50x as my formula and got x=10. K thx @ jygebs

OpenStudy (anonymous):

I did 75+2x=80+1.5x 2x=5+1.5x 0.5x=5 5x=50 x=10 10miles

OpenStudy (anonymous):

K thx.

OpenStudy (anonymous):

yeah 10 miles is correct

OpenStudy (anonymous):

Ty @ Akeller27